# DragFlow: Unleashing DiT Priors with Region Based Supervision for Drag Editing

### v2

@@ Abstract
Drag-based image editing has long suffered from distortions in the target region, largely because the priors of earlier base models, Stable Diffusion, are insufficient to project optimized latents back onto the natural image manifold. With the shift from UNet-based DDPMs to more scalable DiT with flow matching (e.g., SD3.5, FLUX), generative priors have become significantly stronger, enabling advances across diverse editing tasks. However, drag-based editing has yet to benefit from these stronger priors. This work proposes the first framework to effectively harness FLUX’s rich prior for drag-based editing, dubbed DragFlow, achieving substantial gains over baselines. We first show that directly applying point-based drag editing to DiTs performs poorly: unlike the highly compressed features of UNets, DiT features are insufficiently structured to provide reliable guidance for point-wise motion supervision. To overcome this limitation, DragFlow introduces a region-based editing paradigm, where affine transformations enable richer and more consistent feature supervision. Additionally, we integrate pretrained open-domain personalization adapters (e.g., IP-Adapter) to enhance subject consistency, while preserving background fidelity through gradient mask-based hard constraints. Multimodal large language models (MLLMs) are further employed to resolve task ambiguities. For evaluation, we curate a novel Region-based Dragging benchmark (ReD Bench) featuring region-level dragging instructions. Extensive experiments on DragBench-DR and ReD Bench show that DragFlow surpasses both point-based and region-based baselines, setting a new state-of-the-art in drag-based image editing. Code and datasets will be publicly available upon publication.
@@ 1 Introduction
We attribute this limitation to the insufficient generative prior of Stable Diffusion (SD) (Rombach et al., 2022a), the predominant base model, which struggles to constrain optimized latents back onto the natural image manifold. Past findings align with this view: applying nearly identical loss functions for drag editing yields far fewer unnatural distortions when using GAN-based priors compared to SD (Shi et al., 2024b). In parallel, recent advances in generative modeling have shifted from U-Net-based DDPMs to more scalable Diffusion Transformers (DiTs) (Peebles & Xie, 2023) trained with flow matching (Lipman et al., 2022) (e.g., SD 3.5 (Esser et al., 2024a), FLUX.1-dev (Black Forest Labs, 2024)), yielding substantially stronger priors that have propelled progress across various editing tasks (Lu et al., 2025; Yan et al., 2025; Wei et al., 2025; Deng et al., 2024; Wang et al., 2024; Rout et al., 2024). Yet, drag-based editing has not capitalized on these enhanced priors.
@@ 2 Related Work
(i) Optimization-based methods (Xia et al., 2025; Ling et al., 2024; Liu et al., 2024; Zhang et al., 2024c; Jiang et al., 2025; Shi et al., 2024b; Karras et al., 2022; Mou et al., 2023; 2024; Lin et al., 2025; Hou et al., 2024; Cui et al., 2024; Luo et al., 2024; Choi et al., 2024), which is the most prevalent category, iteratively refine inverted noisy latents during inference. These techniques are predominantly point-based: they accept point-wise drag instructions as input and employ motion supervision and point tracking, both of which operate at the point level. However, they often yield unnatural deformations or distortions in the edited images, primarily because the optimized latents deviate from the natural image manifold learned by the base model, residing in out-of-distribution regions. Thus, many studies have focused on more judicious optimization strategies to ensure that the resulting latents can be more readily mapped back to plausible natural images. (ii) Finetuning-based methods (Shin et al., 2024; Shi et al., 2024a), which adapt a base text-to-image (T2I) diffusion model using curated video datasets. Yet, the inherent mismatch between video data and the precise instructions required for drag editing, coupled with the scarcity of high-quality, large-scale training data, limits their generalization. These methods usually fail to achieve complete drag effects and are prone to distortions. (iii) Methods that avoid both finetuning and optimization (e.g., RegionDrag (Lu et al., 2024a) and FastDrag (Zhao et al., 2024f)), instead directly copying and pasting noisy latent patches to target locations computed via predefined mapping functions during inference. While this approach significantly enhances efficiency, it heavily relies on handcrafted priors for the mapping functions, often resulting in edited images that lack faithfulness and realism.
@@ 4.2 Experimental Setup
ReD Bench
RegionDrag (Lu et al., 2024a) NFT 0 1.000 0.957 0.957 33.69 6.38
FastDrag (Zhao et al., 2024f) NFT 0 0.928 0.950 0.941 23.37 5.00
@@
DragBench-DR
RegionDrag (Lu et al., 2024a) NFT 0 1.000 0.942 0.960 32.32 6.31
FastDrag (Zhao et al., 2024f) NFT 0 0.938 0.947 0.952 35.96 6.60
@@ 4.4 Ablation Study
Table 3: Ablation study examining the impact of key components.
@@ Appendix A Additional Related Work
Generative Model-based Image Editing. Recent and significant advancements in generative models (Chang et al., 2023; Ding et al., 2022; Nichol et al., 2021; Ramesh et al., 2022; Rombach et al., 2022a; Saharia et al., 2022; Yu et al., 2022; Peebles & Xie, 2023; Esser et al., 2024a) have enhanced many applications (Avrahami et al., 2023; Ruiz et al., 2023; Hertz et al., 2022; Kim et al., 2022; Tumanyan et al., 2023; Zhou et al., 2025a; b; 2023; 2024a; 2024b; 2024c; Zhao et al., 2024a; d; c; 2025b; 2025a; b; Xie et al., 2024; Wang et al., 2023b; Wang & Chen, 2025a; b; Lu et al., 2024b; c; Li et al., 2025; Ren et al., 2025; Gao et al., 2024; Zhu et al., 2024; 2025; Yu et al., 2025a; b). In this study, we focus on real image editing, which allows users to freely modify actual photographs, producing highly realistic results. Typically, the inputs for image editing include an image and various conditions that help users accurately describe their desired changes. These conditions can encompass text prompts using natural language to specify the edits (Brooks et al., 2023; Zhang et al., 2024b; Fu et al., 2023; Zhang et al., 2024a), region masks to designate areas for modification (Zhao et al., 2024e; Zhuang et al., 2023; Wang et al., 2023a), additional images to provide desired styles or objects (Chen et al., 2024; Lu et al., 2023; Yang et al., 2023), and drag points (Lu & Han, 2025; Nie et al., 2024) that enable users to interactively move specific points in the image to target positions.
@@ B.1 Diffusion Methods
Rectified Flow with ODE Solver
@@ C.5 Details about Experimental Settings
We empirically identify the 17th and 18th double-stream blocks of the DiT backbone as the most effective positions for applying drag optimization. To substantiate this choice, we conduct a visualization-based analysis of the main DiT modules (refer to Fig. 7 and Fig. 8 for details). Our selection is guided by the principle that effective feature blocks should retain high representational fidelity to the input. While certain blocks may encode large amounts of information, much of it can be dominated by noise. By examining PCA visualizations, we qualitatively assess whether the leading components align with the appearance of the original image, thereby confirming that the selected layers capture the essential visual characteristics required for precise and robust editing.
@@
Layers for KV Operations.
Following the ID preservation design introduced in FreeFlux (Wei et al., 2025), we select a subset of layers for KV injection. Specifically, we inject into double-stream blocks at layers $[0,7,8,9,10,18]$ and single-stream blocks at layers $[6,9,18,23,26,31,37]$. During inversion, KV caches are maintained at all valid timesteps, and the corresponding caches are injected back at the corresponding sampling timesteps. To ensure that drag optimization remains consistent with the expected denoising dynamics, we apply the same KV injection during the feature extraction in dragging, consistent with the actual denoising process.
@@ D.3 Leveraging MLLM for Prompt and Tag Generation
The interaction begins with the user providing an input image and roughly indicating the target region via a simple scribble, along with a click specifying the desired target position. The system converts the scribbled region into an initial operational mask representing the affected area. Leveraging an automatic mask generation algorithm Appendix D.2, users do not need to redraw the target region for task specification. The original image and operational mask are then processed by the MLLM, which infers the user’s editing intent and proposes a set of candidate prompts. Specifically, the MLLM generates ten potential prompts paired with corresponding task classes, from which users select the one that best reflects their intended operation.
@@ G.1 Formation of the ReD Benchmark
Figure 11: Examples of real data samples from our ReD benchmark: the first row shows the expected dragging results, where the green region is estimated from the user-specified target centroid (see instruction); the second row presents the source images, while the third row highlights the user-marked operation regions in the form of masks, which may include multiple valid regions; and the last row depicts the adaptively generated masks derived from Appendix D.2. Those two examples correspond to the instructions provided in Appendix G.3.
@@ G.2 Adoption of the DragBench-DR Benchmark
While DragBench-DR extends the benchmark to region-based operations, its evaluation protocol (named MD2) still relies on point-based comparisons: differences are computed between pre-drag and post-drag regions using pre-annotated correspondence points. This design can introduce mismatches, as region-level edits are not faithfully captured by sparse point feature correspondences, leading to potentially unfair assessments of region dragging methods. To better align the evaluation with region-based editing and integrate existing datasets into our experiment, we update the feature comparison criterion by replacing point-based annotations with an automatic centroid-based formulation (see Appendix E.2 for more information).
@@ G.3 Demonstration of Data Samples
We present two real data samples (i.e., SAMPLE (A) and SAMPLE (B)) from our ReD benchmark. The corresponding instructions are provided as follows, and the images are shown in Fig. 11.
@@ Appendix H Extra Qualitative Results
Appendix I LLM Usage Statement

### v3

@@ Abstract
Drag-based image editing has long suffered from distortions in the target region, largely because the priors of earlier base models, Stable Diffusion, are insufficient to project optimized latents back onto the natural image manifold. With the shift from UNet-based DDPMs to more scalable DiT with flow matching (e.g., SD3.5, FLUX), generative priors have become significantly stronger, enabling advances across diverse editing tasks. However, drag-based editing has yet to benefit from these stronger priors. This work introduces DragFlow, the first framework to effectively harness FLUX’s rich prior via region-based supervision, enabling full use of its finer-grained, spatially precise features for drag-based editing and achieving substantial improvements over existing baselines. We first show that directly applying point-based drag editing to DiTs performs poorly: unlike the highly compressed features of UNets, DiT features are insufficiently structured to provide reliable guidance for point-wise motion supervision. To overcome this limitation, DragFlow introduces a region-based editing paradigm, where affine transformations enable richer and more consistent feature supervision. Additionally, we integrate pretrained open-domain personalization adapters (e.g., IP-Adapter) to enhance subject consistency, while preserving background fidelity through gradient mask-based hard constraints. Multimodal large language models (MLLMs) are further employed to resolve task ambiguities. For evaluation, we curate a novel Region-based Dragging benchmark (ReD Bench) featuring region-level dragging instructions. Extensive experiments on DragBench-DR and ReD Bench show that DragFlow surpasses both point-based and region-based baselines, setting a new state-of-the-art in drag-based image editing. Code and dataset are available at https://github.com/Edennnnnnnnnn/DragFlow.
@@ 1 Introduction
We attribute this limitation to the insufficient generative prior of Stable Diffusion (SD) (Rombach et al., 2022a), the predominant base model, which struggles to constrain optimized latents back onto the natural image manifold. Past findings align with this view: applying nearly identical loss functions for drag editing yields far fewer unnatural distortions when using GAN-based priors compared to SD (Shi et al., 2024b). In parallel, recent advances in generative modeling have shifted from U-Net-based DDPMs to more scalable Diffusion Transformers (DiTs) (Peebles and Xie, 2023) trained with flow matching (Lipman et al., 2022) (e.g., SD 3.5 (Esser et al., 2024a), FLUX.1-dev (Black Forest Labs, 2024)), yielding substantially stronger priors that have propelled progress across various editing tasks (Lu et al., 2025; Yan et al., 2025; Wei et al., 2025; Deng et al., 2024; Wang et al., 2024; Rout et al., 2024). Yet, drag-based editing has not capitalized on these enhanced priors.
@@ 2 Related Work
(i) Optimization-based methods (Xia et al., 2025; Ling et al., 2024; Liu et al., 2024; Zhang et al., 2024c; Jiang et al., 2025; Shi et al., 2024b; Karras et al., 2022; Mou et al., 2023; 2024; Lin et al., 2025; Hou et al., 2024; Cui et al., 2024; Luo et al., 2024; Choi et al., 2024), which is the most prevalent category, iteratively refine inverted noisy latents during inference. These techniques are predominantly point-based: they accept point-wise drag instructions as input and employ motion supervision and point tracking, both of which operate at the point level. However, they often yield unnatural deformations or distortions in the edited images, primarily because the optimized latents deviate from the natural image manifold learned by the base model, residing in out-of-distribution regions. Thus, many studies have focused on more judicious optimization strategies to ensure that the resulting latents can be more readily mapped back to plausible natural images. (ii) Finetuning-based methods (Shin et al., 2024; Shi et al., 2024a), which adapt a base text-to-image (T2I) diffusion model using curated video datasets. Yet, the inherent mismatch between video data and the precise instructions required for drag editing, coupled with the scarcity of high-quality, large-scale training data, limits their generalization. These methods usually fail to achieve complete drag effects and are prone to distortions. (iii) Methods that avoid both finetuning and optimization (e.g., RegionDrag (Lu et al., 2024a) and FastDrag (Zhao et al., 2024d)), instead directly copying and pasting noisy latent patches to target locations computed via predefined mapping functions during inference. While this approach significantly enhances efficiency, it heavily relies on handcrafted priors for the mapping functions, often resulting in edited images that lack faithfulness and realism.
@@ 4.2 Experimental Setup
ReD Bench RegionDrag (Lu et al., 2024a) NFT 0 1.000 0.957 0.957 33.69 6.38
FastDrag (Zhao et al., 2024d) NFT 0 0.928 0.950 0.941 23.37 5.00
@@
DragBench-DR RegionDrag (Lu et al., 2024a) NFT 0 1.000 0.942 0.960 32.32 6.31
FastDrag (Zhao et al., 2024d) NFT 0 0.938 0.947 0.952 35.96 6.60
@@ 4.4 Ablation Study
Table 3: Ablation study on ReD Bench examining the impact of some key components.
@@ 5 Conclusion
Acknowledgments
This research is supported by the National Research Foundation, Prime Minister’s Office, Singapore, and the Ministry of Digital Development and Information, under its Online Trust and Safety (OTS) Research Programme (MDDI-OTS-001). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Prime Minister’s Office, Singapore, the Ministry of Digital Development and Information, or the Centre for Advanced Technologies in Online Safety.
@@ Appendix A Additional Related Work
Generative Model-based Image Editing.
Recent and significant advancements in generative models (Chang et al., 2023; Ding et al., 2022; Nichol et al., 2021; Ramesh et al., 2022; Rombach et al., 2022a; Saharia et al., 2022; Yu et al., 2022; Peebles and Xie, 2023; Esser et al., 2024a) have enhanced many applications (Avrahami et al., 2023; Ruiz et al., 2023; Hertz et al., 2022; Kim et al., 2022; Tumanyan et al., 2023; Zhou et al., 2025; 2023; 2024a; 2024b; 2024c; Zhao et al., 2024a; 2025b; 2025a; b; 2026; Wang et al., 2025; Wang and Chen, 2025a; b; Lu et al., 2024b; c; Li et al., 2025; Ren et al., 2025; Gao et al., 2024; 2025; Zhu et al., 2024; 2025; Yu et al., 2025a; b). In this study, we focus on real image editing, which allows users to freely modify actual photographs, producing highly realistic results. Typically, the inputs for image editing include an image and various conditions that help users accurately describe their desired changes. These conditions can encompass text prompts using natural language to specify the edits (Brooks et al., 2023; Zhang et al., 2024b; Fu et al., 2023; Zhang et al., 2024a), region masks to designate areas for modification (Zhao et al., 2024c; Zhuang et al., 2023; Wang et al., 2023), additional images to provide desired styles or objects (Chen et al., 2024; Lu et al., 2023; Yang et al., 2023), and drag points (Lu and Han, 2025; Nie et al., 2024) that enable users to interactively move specific points in the image to target positions.
@@ B.1 Diffusion Methods
Rectified Flow with ODE Solver.
@@ C.5 Details about Experimental Settings
We empirically identify the 17th and 18th double-stream blocks of the DiT backbone as the most representative features to facilitate drag optimization. To substantiate this choice, we conduct a visualization-based analysis of the main DiT modules (refer to Fig. 7 and Fig. 8 for details). Our selection is guided by the principle that effective feature blocks should retain high representational fidelity to the input. While certain blocks may encode large amounts of information, much of it can be dominated by noise. By examining PCA visualizations, we qualitatively assess whether the leading components align with the appearance of the original image, thereby confirming that the selected layers capture the essential visual characteristics required for precise and robust editing.
@@
Figure 9: Comparison of dragging effectiveness across feature manipulation layers in the FLUX model. The analysis is evaluated on ReD Bench, demonstrating that double-stream layers (prefixed with “D”, in blue) consistently outperform single-stream layers (prefixed with “S”, in purple).
Layer-based Performance Comparison.
To further validate the effectiveness of our selected feature manipulation layers, we conducted an extensive ablation study focused on layer-wise feature selection. Specifically, under the same configuration, we compute the loss each time using the features from a particular layer. This procedure was repeated across all 57 layers from both the double-stream and single-stream branches of the FLUX model, enabling a systematic comparison of editing behaviors across various layers.
The results are summarized in Fig. 9. We evaluate three metrics of dragging effectiveness: MD1, IFs2s, and IFs2t. As shown, operations performed on the double-stream layers (i.e., layers prefixed with “D”) consistently outperform those on the single-stream layers (i.e., layers prefixed with “S”) under the same dragging configuration. Notably, the 17th and 18th double-stream blocks (i.e., D17 and D18), which we ultimately adopt in our framework, exhibit the optimal performance robustly across all these metrics. This indicates that our chosen layers yield the highest-effectiveness solutions in terms of spatial displacement precision, preservation of post-dragging local features, and effective suppression of original-location features.
@@
C.6 Quantitative Analysis for DiT Supervision Granularity
Table 4: Quantitative comparison on DragBench-DR between Point-based and Region-based strategies for the DiT architecture. Both methods utilize the same backbone and auxiliary modules.
Attempts Image Fidelity Mean Distance
Point-based variant 0.961 0.943 0.958 37.02 6.94
Region-based variant (DragFlow) 0.969 0.948 0.941 31.59 5.93
To further validate the analysis in Sec. 3.1, that DiTs’ fine-grained features render sparse point supervision ineffective, we conducted a controlled quantitative comparison.
By evaluating two distinct supervision paradigms: a baseline Point-based variant (i.e., applying standard point to the DiT features) and our proposed Region-based variant (DragFlow). To ensure a rigorous evaluation of the supervision mechanisms in isolation, both variants were equipped with identical auxiliary modules, specifically Background Preservation and Adapter-Enhanced Inversion.
The comparative results are summarized in Tab. 4. These quantitative findings empirically corroborate our assertion that sparse point tracking is insufficient for the high-frequency feature landscape of DiTs, whereas region-level affine supervision effectively bridges this gap to fully unleash the model’s generative potential.
@@ D.3 Leveraging MLLM for Prompt and Tag Generation
The interaction begins with the user providing an input image and roughly indicating the target region via a simple scribble, along with a click specifying the desired target position. The system converts the scribbled region into an initial operational mask representing the affected area. Leveraging an automatic mask generation algorithm (Appendix D.2), users do not need to redraw the target region for task specification. The original image and operational mask are then processed by the MLLM, which infers the user’s editing intent and proposes a set of candidate prompts. Specifically, the MLLM generates ten potential prompts paired with corresponding task classes, from which users select the one that best reflects their intended operation.
@@
D.4 Effectiveness of MLLM-driven Intent Parsing
To quantitatively assess the contribution and robustness of the MLLM-driven intent parsing module, we conducted an extensive ablation study on the DragBench-DR dataset. This analysis compares the performance of DragFlow under three distinct experimental scenarios:
  • (1) Null Prompt: The optimization is driven solely by image and mask inputs, removing the text guidance entirely;
  • (2) Incorrect Prompt: The model is provided with intentionally misinterpreted prompts (e.g., describing a rotation task as deformation) to simulate worst-case MLLM failure;
  • (3) Matched Prompt: The standard configuration using prompts generated by MLLMs (comparing both GPT-5 and QWen-VL) based on the image and operation inputs.
The averaged quantitative results are presented in Tab. 5. As evidenced by the data, the framework remains robust even without textual guidance (i.e., Null Prompt), demonstrating the efficacy of the core region-level affine supervision.
In contrast, including Matched Prompts provides a valuable performance boost, whereas Incorrect Prompts degrade results due to conflicting guidance. However, thanks to the “Double Confirmation” workflow (detailed in Appendix D.3) we implemented, it helps prevent such misinterpretations in practice: rather than autonomous execution, the system requires users to select the optimal intent from generated candidates before optimization, ensuring the process is driven strictly by accurate, user-validated instructions.
Table 5: To validate the effectiveness of prompts automatically generated by MLLM, we conducted an additional test. First, we examined whether using accurate and matching prompts is effective for drag operations. We considered three scenarios: Null Prompt, Incorrect Prompt (i.e., intentionally misinterpreted prompts), and Matched Prompt generated by MLLMs based on image and operation inputs. For the last scenario, we examine the effectiveness of two MLLMs (GPT-5 vs. QWen-VL). This experiment was conducted and scored on DragBench-DR.
DragFlow w/ Image Fidelity Mean Distance
Null Prompt 0.966 0.944 0.948 33.81 6.81
Incorrect Prompt 0.955 0.936 0.955 36.87 7.73
Matched Prompt (by QWem-VL) 0.968 0.947 0.945 32.42 6.25
Matched Prompt (by GPT-5) 0.969 0.948 0.941 31.59 5.93
@@ Appendix F Additional Baseline Information
And some encapsulated modules applied in our framework or experiments:
  • 1
    InstantCharacter: https://github.com/Tencent-Hunyuan/InstantCharacter
  • 2
    SD3.5-Large-IP-Adapter: https://huggingface.co/InstantX/SD3.5-Large-IP-Adapter
@@ G.1 Formation of the ReD Benchmark
Figure 12: Real samples from ReD Bench: the first colume shows the dragging preview, where the green region is estimated from the user-specified target centroid; the second colume presents the source images, while the third colume highlights the user-marked operation regions in the form of masks, which may include multiple valid regions; and the last colume depicts the adaptive masks generated (detailed in Appendix D.2). Other matched instructions are provided in Appendix G.3.
@@ G.2 Adoption of the DragBench-DR Benchmark
While DragBench-DR extends the benchmark to region-based operations, its evaluation protocol (MD2) still relies on point comparisons: differences are computed between pre- and post-drag regions using pre-annotated correspondence points. This design can introduce mismatches, as region-level edits are not faithfully captured by sparse point feature correspondences, leading to potentially unfair assessments for region methods. To better align the evaluation with region-based editing and integrate existing datasets into our experiment, we update the feature comparison criterion by replacing point-based annotations with an automatic centroid-based formulation (see Appendix E.2).
@@ G.3 Demonstration of Data Samples
We present two real data samples (i.e., SAMPLE (A) and SAMPLE (B)) from the ReD Bench. The corresponding instructions are provided as follows, and the images are shown in Fig. 12.
@@ H.1 Additional Qualitative Samples
H.2 Challenging Scenarios
To further assess the robustness of DragFlow, we evaluate its performance under extreme conditions. As illustrated in Fig. 14, our analysis focuses on two distinct dimensions: (a) complex operational requirements and (b) unusual feature structures. Following this, in Subsec. H.3, we are also advised to provide some illustrations about the failure cases. We believe these parts may inspire and offer insights for future research about image drag-editing.
(a) Cases with Complex Instructions.
Fundamentally, we conceptualize the defined operations as elementary “building blocks.” Consequently, sophisticated instructions that involve composite movements (e.g., do both rotation and relocation) or non-affine warps (e.g., bending) are effectively executed by composing these basic units. Among the composite movements, we consider two scenarios, where multiple distinct operations are processed in parallel or in sequence. The lower section of Fig. 14 demonstrates how our framework effectively handles these complex instructions and overcomes the challenging scenarios.
(b) Cases with Complex Structures.
Regarding the non-rigid deformations inherent to complex textures such as hair and cloth, our empirical results confirm that the affine assumption remains valid. By applying flexible single operations or iteratively composing multiple blocks to approximate the target motion, the framework maintains high editing effectiveness and structural fidelity even on these intricate features. Representative samples validating this capability are presented in the upper part of Fig. 14.
H.3 Failure Scenarios
Figure 13: Extra qualitative results for failure scenarios. Left: The incorrect operation mask and failure outcomes; Right: The suitable operation mask with failure outcomes.
Dependency on Input Precision.
A primary limitation of the proposed framework is its reliance on the quality of user-provided inputs. Given that both the region-level affine supervision and the adaptive gradient masking mechanisms hinge directly on the initial mask annotation, ambiguous or inaccurate inputs can inevitably introduce generation artifacts.
Empirically, we observe that while the framework exhibits considerable tolerance for slightly oversized masks, ensuring complete coverage of the targeted semantic object is critical. Specifically, undersized or incomplete masks tend to compromise the editing quality. Fig. 13 illustrates this phenomenon by comparing outcomes derived from precise versus careless annotations. To relax these strict precision requirements, future iterations could explore integrating auxiliary techniques such as superpixel segmentation or implementing error-tolerant peripheral buffer zones to robustly handle manual variances.
Appendix I Runtime and Memory Comparison
To provide a comprehensive understanding of DragFlow’s practical performance, especially as an optimization-based method, a quantitative analysis of its computational cost is necessary. This discussion is intended to help users understand the quality versus speed trade-offs inherent in the approach. Tab. 6 presents a detailed comparison of the average inference time and GPU memory consumption for DragFlow alongside various baseline methods. The results were recorded as an average over all samples contained in DragBench-DR using one NVIDIA H100 GPU.
In our implementation, we applied qint8 quantization to the FLUX backbone to accelerate inference and reduce memory usage during the optimization. Additionally, by enabling CPU Offloading, the framework is capable of running on a consumer-grade $24$ GB GPU. Beyond the current level, we believe there is still plenty of room to further reduce memory costs and increase speed through future improvements in both the algorithm and the code.
As the data illustrates, DragFlow’s resource requirements are higher than the baseline methods. This is an anticipated result and a direct consequence of the underlying model architecture. DragFlow is the only method in this comparison engineered to operate on the FLUX.1 model, a high-parameter, DiT-based architecture. In contrast, all other baselines are built upon the significantly smaller and less complex Stable Diffusion models.
Table 6: This table collects and summarizes the inference time and memory consumption depending on different image drag-editing methods. Among them, DragFlow, as the only DiT-based dragger, shows the highest resource demand due to its more complex backbone architecture. The results are recorded as averages on DragBench-DR.
Methods Prep + Edit Time (s) Peak Memory (GB)
RegionDrag $9.2$ $4.3$
FastDrag $5.4$ $5.2$
InstantDrag $1.2$ $4.1$
DragLoRA $58.6$ $14.6$
FreeDrag $113.2$ $13.2$
DragNoise $103.4$ $12.2$
GoodDrag $104.6$ $12.8$
CLIPDrag $101.0$ $19.4$
DragDiffusion $82.1$ $12.1$
DragFlow $158.7$ $23.5$
Appendix J Framework and Module Generalizability
Table 7: The editing performance of our DragFlow framework on different DiT-based generative priors (SD3.5 vs. FLUX.1). The results are averaged on DragBench-DR.
Method Image Fidelity Mean Distance
DragFlow (based on Stable Diffusion 3.5) 0.962 0.945 0.952 35.21 6.58
DragFlow (based on FLUX.1-dev) 0.969 0.948 0.941 31.59 5.93
Table 8: The editing performance of our DragFlow framework on different ID-preservation modules (IP-Adapter vs. LoRA). The results are averaged on DragBench-DR.
Method Image Fidelity Mean Distance
DragFlow (w/ LoRA) 0.970 0.949 0.940 30.98 5.88
DragFlow (w/ IP-Adapter) 0.969 0.948 0.941 31.59 5.93
J.1 Generalizability of Dragging Framework
To validate the generalization capability of DragFlow beyond the specific architecture of FLUX, we conducted an additional evaluation applying our framework to Stable Diffusion 3.5 (SD3.5), another prominent text-to-image model based on the DiT architecture. Notably, since the default InstantCharacter adapter is FLUX-specific, we employed SD3.5-Large-IP-Adapter to ensure a fair comparison. Tab. 7 presents the comparative editing performance averaged over all samples from the DragBench-DR dataset. As the results indicate, DragFlow maintains robust performance when transferred to the SD3.5 backbone.
Specifically, the SD3.5-based implementation achieves an MD1 of $35.21$ and an IFs2t of $0.945$, versus $31.59$ and $0.948$ of the FLUX edition, respectively. The FLUX-based version demonstrates a slight performance advantage, likely due to its substantially larger parameter count and enhanced generative prior, while the performance on SD3.5 remains competitive and consistent. These findings provide empirical evidence that the key contributions of DragFlow, including region-level affine supervision, gradient mask constraints, and adapter-enhanced inversion, are agnostic to the architecture used. The framework effectively generalizes across different DiT-based backbones, confirming that its efficacy is not limited to the specific training paradigm of FLUX but rather serves as a generalized solution for modern DiT generative models.
J.2 Generalizability of ID-preservation Design
To further evaluate the design choices behind our method, we conducted an experiment in which the IP-Adapter was replaced with a subject-specific LoRA. In this setup, a LoRA was first trained for the target subject and then applied during the drag-editing stage. Performance on DragBench-DR is summarized in Table 8.
The LoRA-based variant exhibits slightly stronger performance, which aligns with expectations since the LoRA is expressly optimized for the specific subject rather than functioning as a generic, open-domain adapter. However, this improvement comes with the overhead of training a new LoRA for each subject. In contrast, many modern text-to-image models now include high-quality, readily usable IP-Adapters, making the IP-Adapter approach more practical and broadly applicable in typical workflows.
Appendix K LLM Usage Statement